# Order-Optimal Instance-Dependent Bounds for Offline Reinforcement Learning with Preference Feedback

### v1

@@ Appendix E Proof of Upper bound
Appendix F Theoretical analysis of DP-RL-Low
Lemma F.1.
@@ Appendix F Theoretical analysis of DP-RL-Low
Proof of Lemma F.1.
For simplicity of notation, we let $x_{0}\coloneqq\mathbb{E}(\tilde{X}_{n})$, which implies $x_{0}=p\in[1-\beta,\beta]$. Again, by the smoothness of $f(\cdot)$ on $[1-\beta,\beta]$, we get the Taylor expansion of $f(\cdot)$ on $[1-\beta,\beta]$,
$f(x)=\begin{cases}f(x_{0})+f^{\prime}(x_{0})\cdot(x-x_{0})+\frac{1}{2}f^{\prime\prime}(\xi_{x})\cdot(x-x_{0})^{2}\quad\text{ if }x\in[1-\beta,\beta]\\ f(x_{0})+f^{\prime}(x_{0})\cdot(\beta-x_{0})+\frac{1}{2}f^{\prime\prime}(\xi_{\beta})\cdot(\beta-x_{0})^{2}\quad\text{ if }x>\beta\\ f(x_{0})+f^{\prime}(x_{0})\cdot(1-\beta-x_{0})+\frac{1}{2}f^{\prime\prime}(\xi_{\beta})\cdot(1-\beta-x_{0})^{2}\quad\text{ if }x<1-\beta,\\ \end{cases}$ (72)
@@
where ${\rm Var}(\cdot)$ represents the variance and $[x]_{+}=\max\{x,0\}$.
@@
Then, by Lemma A.5, we get that
@@
Similarly, from Lemma A.5, we further get that
@@
Note that by definition, we have $\|f(\tilde{X}_{n})-f(p)\|_{\rm vp}^{2}=\|f(\tilde{X}_{n})\|_{\rm vp}^{2}$, which completes the proof of Lemma E.2. ∎
@@
Fix any consistent instance $v$ and $n>0$ under DP-RL-LOW. By lemma E.2 and lemma A.3, we get that $\tilde{r}_{k,i^{*}_{k},i}$ is subGaussian with variance proxy to be,
@@
In addition, by Lemma F.1, we get that for any $(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}$ with $N_{k^{\prime},i^{\prime},j^{\prime}}>0$,
@@
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathbbm{1}_{\{\tilde{r}_{k,i^{*}_{k},i}<0\}}\cdot\rho_{k}\Delta_{k,i}$ (84)
$\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\bigg{[}\big{(}\mathbbm{1}_{\{\tilde{r}_{k,i^{*}_{k},i}\leq\mathbb{E}_{v}^{\rm DP\mbox{-}RL\mbox{-}LOW}\left[\tilde{r}_{k,i^{*}_{k},i}\right]-\Delta_{k,i}/2\}}$
$\hskip 85.35826pt\lor\mathbbm{1}_{\{\lvert\mathbb{E}_{v}^{\rm DP\mbox{-}RL\mbox{-}LOW}\left[\tilde{r}_{k,i^{*}_{k},i}\right]-r_{k,i^{*}_{k},i}\rvert\leq\Delta_{k,i}/2\}}\big{)}\cdot\rho_{k}\Delta_{k,i}\bigg{]}$ (85)
By Lemma A.4 and  (79), we get that for any $(k,i)\in\mathcal{S}\times\mathcal{A}$ with $i\neq i^{*}_{k}$,

### v2

@@ Appendix H Proof of Upper bounds
Proof of Theorem 5.1.
Fix any consistent instance $v$ under RL-LOW-MDP. By Lemma H.2 and Lemma D.3, we get that $\hat{r}_{k,i,j}$ is subGaussian with variance proxy as
$\|\hat{r}_{k,i,j}\|_{\rm vp}^{2}\leq\frac{C\gamma_{k,i,j}}{n},$ (75)
where $C\leq(6\sqrt{2e}\cdot(3\sqrt{\log 2}+1))^{2}\cdot\frac{3}{2(1-\beta)^{4}}$ and $\beta=\frac{\exp(2L)}{1+\exp(2L)}$, and $\gamma_{k,i,j}$ is defined as
$\gamma_{k,i,j}\coloneqq\sum_{k^{\prime}\in\mathcal{S},i^{\prime},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{(w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}})^{2}}{N_{k^{\prime},i^{\prime},j^{\prime}}}$
In addition, by Lemma H.1, we get that for any $(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{A}$ with $N_{k^{\prime},i^{\prime},j^{\prime}}>0$,
$\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[f(B_{k^{\prime},i^{\prime},j^{\prime}})\right]-f(\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[B_{k^{\prime},i^{\prime},j^{\prime}}\right])\rvert\leq\frac{3}{(1-\beta)^{4}\sqrt{n\cdot N_{k^{\prime},i^{\prime},j^{\prime}}}},$
which implies for any $(k,i,j)\in\mathcal{S}\times\mathcal{A}^{2}$ with $i\neq j$,
$\left\lvert\sum_{k^{\prime}\in\mathcal{S},i^{\prime}\in\mathcal{A},j^{\prime}\in\mathcal{A}}w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}}\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[f(B_{k^{\prime},i^{\prime},j^{\prime}})\right]-f(\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[\hat{r}_{k,i,j}\right])\right\rvert$
$\leq\sum_{k^{\prime}\in\mathcal{S},i^{\prime}\in\mathcal{A},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{3\lvert w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}}\rvert}{(1-\beta)^{4}\sqrt{n\cdot N_{k^{\prime},i^{\prime},j^{\prime}}}}.$ (76)
That is,
$\left\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[\hat{r}_{k,i,j}\right]-r_{k,i,j}\right\rvert$ $\leq\sum_{k^{\prime}\in\mathcal{S},i^{\prime}\in\mathcal{A},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{3\lvert w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}}\rvert}{(1-\beta)^{4}\sqrt{n\cdot N_{k^{\prime},i^{\prime},j^{\prime}}}}$ (77)
$=\frac{3\tilde{\gamma}_{k,i,j}}{(1-\beta)^{4}\sqrt{n}},$ (78)
where we denote
$\tilde{\gamma}_{k,i,j}\coloneqq\sum_{k^{\prime}\in\mathcal{S},i^{\prime}\in\mathcal{A},j^{\prime}\in\mathcal{A}:N_{k^{\prime},i^{\prime},j^{\prime}}>0}\frac{\lvert w^{(k,i,j)}_{k^{\prime},i^{\prime},j^{\prime}}\rvert}{\sqrt{N_{k^{\prime},i^{\prime},j^{\prime}}}}$
and recall that $r_{k,i,j}\coloneqq r_{k,i}-r_{k,j}$. In addition, we denote $r(\pi)\coloneqq\mathbb{E}_{k\sim d^{\pi}}[r_{k,\pi(k)}]$ for any MDP policy $\pi$.
Note that by the definition of $\hat{\pi}_{\rm out}$ under RL-LOW-MDP, we have
$R^{\rm MDP}(\hat{\pi}_{\rm out})$ $\leq\sum_{\pi\neq\hat{\pi}_{\rm out}}\mathds{1}_{\{r(\hat{\pi}_{\rm out})<r(\pi)\}}\cdot R^{\rm MDP}(\pi)$
$\leq\sum_{\pi\neq\pi^{*}}R^{\rm MDP}(\pi)\cdot\mathds{1}_{\{\bigcup_{k\in\mathcal{S}}\{\hat{r}_{k,\pi(k),\pi^{*}(k))}-r_{k,\pi(k),\pi^{*}(k)}\geq R^{\rm MDP}(\pi)\}\}}$ (79)
By Lemma D.4 and Eqn. (75), we get that for any $\pi\neq\pi^{*}$ and $k\in\mathcal{S}$ with $\pi(k)\neq\pi^{*}(k)$,
$\mathbb{P}_{v}^{\rm RL-LOW-MDP}\bigg(\hat{r}_{k,\pi(k),\pi^{*}(k)}\geq\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[\hat{r}_{k,\pi(k),\pi^{*}(k)}\right]+R^{\rm MDP}(\pi)/2\bigg)$
$\leq\exp\left(-\frac{2n(R^{\rm MDP}(\pi))^{2}}{C\gamma_{k,\pi(k),\pi^{*}(k)}}\right),$ (80)
and by Eqn. (78) we get that
$\big\lvert\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[\hat{r}_{k,i,j}\right]-r_{k,i,j}\big\rvert\leq R^{\rm MDP}(\pi)/2$
for all $n>\frac{18\tilde{\gamma}_{k,i,j}^{2}}{(1-\beta)^{8}R^{\rm MDP}(\pi)/2}$. That is, for all sufficiently large $n$, we have
$\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[R^{\rm MDP}(\hat{\pi}_{\rm out})\right]\leq\sum_{\pi\neq\pi^{*}}R^{\rm MDP}(\pi)\sum_{k\in\mathcal{S}:\pi(k)\neq\pi^{*}(k)}\exp\left(-\frac{2n(R^{\rm MDP}(\pi))^{2}}{C\gamma_{k,\pi(k),\pi^{*}(k)}}\right)$
which further implies that for all sufficiently large $n$, we have
$\mathbb{E}_{v}^{\rm RL\mbox{-}LOW\mbox{-}MDP}\left[R^{\rm MDP}(\hat{\pi}_{\rm out})\right]\leq\exp\left(-\frac{n}{C_{\rm MDP}\cdot H^{\rm MDP}(v)}\right).$
This completes the proof of Theorem 5.1.
Appendix I Theoretical analysis of DP-RL-Low
I.1 instance-dependent upper bound
Lemma I.1.
@@ Appendix I Theoretical analysis of DP-RL-Low
Proof of Lemma I.1.
For simplicity of notation, we let $x_{0}\coloneqq\mathbb{E}(\tilde{X}_{n})$, which implies $x_{0}=p\in[1-\beta,\beta]$. Again, by the smoothness of $f(\cdot)$ on $1-\beta,\beta]$, we get the Taylor expansion of $f(\cdot)$ on $1-\beta,\beta]$,
[…]
@@
where ${\rm Var}(\cdot)$ represents the variance and $x]_{+}=\max\{x,0\}$.
@@
Then, by Lemma D.5, we get that
@@
Similarly, from Lemma D.5, we further get that
@@
Note that by definition, we have $\|f(\tilde{X}_{n})-f(p)\|_{\rm vp}^{2}=\|f(\tilde{X}_{n})\|_{\rm vp}^{2}$, which completes the proof of Lemma H.2. ∎
@@
Fix any consistent instance $v$ and $n>0$ under DP-RL-LOW. By lemma H.2 and lemma D.3, we get that $\tilde{r}_{k,i^{*}_{k},i}$ is subGaussian with variance proxy to be,
@@
In addition, by Lemma I.1, we get that for any $(k^{\prime},i^{\prime},j^{\prime})\in\mathcal{S}\times\mathcal{A}^{2}$ with $N_{k^{\prime},i^{\prime},j^{\prime}}>0$,
@@
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathds{1}_{\{\tilde{r}_{k,i^{*}_{k},i}<0\}}\cdot\rho_{k}\Delta_{k,i}$ (93)
$\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\bigg[\big(\mathds{1}_{\{\tilde{r}_{k,i^{*}_{k},i}\leq\mathbb{E}_{v}^{\rm DP\mbox{-}RL\mbox{-}LOW}\left[\tilde{r}_{k,i^{*}_{k},i}\right]-\Delta_{k,i}/2\}}$
$\hskip 85.35826pt\lor\mathds{1}_{\{\lvert\mathbb{E}_{v}^{\rm DP\mbox{-}RL\mbox{-}LOW}\left[\tilde{r}_{k,i^{*}_{k},i}\right]-r_{k,i^{*}_{k},i}\rvert\leq\Delta_{k,i}/2\}}\big)\cdot\rho_{k}\Delta_{k,i}\bigg]$ (94)
By Lemma D.4 and  (88), we get that for any $(k,i)\in\mathcal{S}\times\mathcal{A}$ with $i\neq i^{*}_{k}$,
@@
I.2 Worst-case Upper Bound
[…]
Proposition I.3.
(Worst-Case Upper Bound for DP-RL-LOW) For any consistent instance $v$ and for all $n\geq 1$,
$\mathbb{E}^{\rm DP-RL\mbox{-}LOW}_{v}\left[R_{n}\right]\!\leq\!C_{\rm WDP}\cdot\left(\frac{\sum\limits_{\begin{subarray}{c}k,i:i\neq i^{*}_{k}\end{subarray}}\rho_{k}(\sqrt{\gamma_{k,i}}\!+\!\tilde{\gamma}_{k,i})}{\sqrt{n}}+\frac{\sum\limits_{\begin{subarray}{c}k,i:i\neq i^{*}_{k}\end{subarray}}\rho_{k}(\sqrt{\gamma_{k,i}^{\rm DP}}\!+\!\tilde{\gamma}^{\rm DP}_{k,i})\sqrt{\log(1.25/\delta)}}{\epsilon{n}}\right)$ (96)
where $C_{\rm WDP}>0$ is a universal constant.
Proof of Proposition I.3.
Note that by the definition of $R_{n}$ we have
$R_{n}$ $\leq\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\mathds{1}_{\{\tilde{r}_{k,i^{*}_{k},i}\leq 0\}}\cdot\rho_{k}\Delta_{k,i}$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\rho_{k}\bigg[\mathds{1}_{\{\Delta_{k,i}<\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}+\frac{48\tilde{\gamma}^{\rm DP}_{k,i}\sqrt{{{\log(1.25/\delta)}}}}{(1-\beta)^{4}{{(\varepsilon n)}}}\}}\cdot\big(\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}+\frac{8\tilde{\gamma}^{DP}_{k,i}\sqrt{{{\log(1.25/\delta)}}}}{(1-\beta)^{4}{{(\varepsilon n)}}}\big)$
$\hskip 85.35826pt+\mathds{1}_{\{\lvert\mathbb{E}_{v}^{\rm DP-RL\mbox{-}LOW}\left[\tilde{r}_{k,i,i^{*}_{k}}\right]-r_{k,i,i^{*}_{k}}\rvert\leq\Delta_{k,i}/2\}}\cdot\Delta_{k,i}\bigg]$ (97)
Hence, by Eqn. (97) and (I.1), we have
$\mathbb{E}_{v}^{\rm DP-RL\mbox{-}LOW}[R_{n}]$ ${\leq}\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\rho_{k}\bigg[\big(\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}+\frac{8\tilde{\gamma}^{DP}_{k,i}\sqrt{{{\log(1.25/\delta)}}}}{(1-\beta)^{4}{{(\varepsilon n)}}}\big)$
$\hskip 21.68121pt+\exp\left(-\frac{2\Delta_{k,i}^{2}}{C\cdot(\gamma_{k,i}/n+{8\gamma_{k,i}^{\rm DP}\log(1.25/\delta)}/{(\varepsilon n)^{2}})}\right)\Delta_{k,i}\bigg]$
${\leq}\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\rho_{k}\bigg[\big(\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}\sqrt{n}}+\frac{8\tilde{\gamma}^{DP}_{k,i}\sqrt{{{\log(1.25/\delta)}}}}{(1-\beta)^{4}{{(\varepsilon n)}}}\big)$
$\hskip 21.68121pt+\sqrt{\frac{C\cdot\gamma_{k,i}}{2n}}+\sqrt{\frac{C\cdot{8\gamma_{k,i}^{\rm DP}\log(1.25/\delta)}}{2(\varepsilon n)^{2}}}\bigg]$
${=}\sum_{(k,i)\in\mathcal{S}\times\mathcal{A}:i\neq i^{*}_{k}}\rho_{k}\bigg[\frac{1}{\sqrt{n}}\cdot\left(\frac{6\tilde{\gamma}_{k,i}}{(1-\beta)^{4}}+\sqrt{\frac{C\cdot\gamma_{k,i}}{2}}\right)$
$\hskip 21.68121pt+\frac{1}{\varepsilon n}\cdot\left(\frac{8\tilde{\gamma}^{\rm DP}_{k,i}\sqrt{{{\log(1.25/\delta)}}}}{(1-\beta)^{4}{}}+\sqrt{\frac{C\cdot{8\gamma_{k,i}^{\rm DP}\log(1.25/\delta)}}{2}}\right)\bigg]$
This completes the desired proof. ∎